Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Sat Nov 8 14:00:32 MET 1997) Success

Medicina (venturi at Sat Nov 8 14:18:17 MET 1997) Success
rain inthe night

Westerbork (tony foley at Sat Nov 8 14:51:53 MET 1997) Minor failures
Total power dropouts from ~2319UT to 0030UT. Cause unknown, but seems to be happening in the frontend. Still no phasecal.

Jodrell (Alastair Gunn at Sat Nov 8 17:43:59 MET 1997) Minor failures
IF attenuation changed at start of experiment. 5-10 minutes of data probably lost at around 16:15 due to telescope control failure; source command resent. Ignore errors about track interface not in remote. Telescope azimuth encoder failure at about 06:03 but fault cleared quickly. Telescope on source from about 06:06 onwards. Weather overcast. Track 21 dead.

Onsala (A. Polatidis at Sat Nov 8 19:06:32 MET 1997) Minor failures
Guess what ? Strong winds forced us to stow the telescope between 2040 and 0340 UT. Otherwise no other known problems. Tapes were left to run so no data on second tape after 2040 and data begins on third tape after 0345.

HartRAO (Jonathan Quick at Mon Nov 10 08:55:40 MET 1997) Success
No problems apert from some ripple on IF

Effelsberg (dgraham at Mon Nov 10 09:59:37 MET 1997) Minor failures
missed 311 1944 to 2025 tape problem. Some rain around 312 0800

Noto (C.Trigilio at Fri Nov 14 09:28:06 MET 1997) Success

Urumqi (Li Suqin at Fri Nov 14 11:16:10 MET 1997) Minor failures
It is snow. H-mase to GPS =2.87us,Format to GPS =2.80us. From 20:20:00 to 21:06:04 not recorded data.

Calibration (Fredrik T Rantakyro at Wed Dec 3 16:40:40 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name eh003.antab Read the comments in the file to get further information.